OpenStudy (anonymous):

A battery produces 40.8V when 7.10 A are drawn from it and 47.3V when 2.20 A are drawn. What is the emf and internal resistance of the battery?

OpenStudy (anonymous):

not sure how to solve this without knowing the open circuit voltage(ocv). A battery's IR is equal to the difference between the OCV and the CCV, divided by the applied load. This formula is derived from ohm's law E/IxR. For example: If the OCV=14V and the CCV=12V with a 100A load applied, the internal resistance is 0.02ohms. 14-12=2 2/100=.02. Without having the OCV, i aint sure how to do it. sorry.

OpenStudy (stormfire1):

You need to write out the two equations with two unknowns and solve it: E - 7.2 R = 40.8V E - 2.5 R = 47.0V E =~ 50.3 and the internal resistance =~ 1.32 ohms